# Large Language Models for Scholarly Ontology Generation: An Extensive Analysis in the Engineering Field

### v1

@@ 5.2 Standard prompting, two-way strategy
Among the full open models, mixtral achieves an average F1-score of 0.847, significantly outperforming both llama-2 and mistral. It is able to identify the broader, narrower, and same-as with F1-score around 0.82, but shows exceptional capabilities (0.925) in determining the relationship other.
Regarding the quantised models, there are three models (e.g., dolphin-mistral-dpo, dolphin-openorca, and openchat) that obtain high values of average F1-score above 0.78.
@@
In this case, the full open models and their quantised counterparts exhibit similar performance levels (e.g., 0.847 for mixtral and 0.853 for dolphin-openorca). Moreover, the two-way strategy has effectively reduced the performance gap between these models and their larger, proprietary counterparts, narrowing the difference to just 0.113 F1 points.
@@ 5.3 Chain-of-thought prompting, one-way strategy
Table 3 presents the results of the chain-of-thought, one-way strategy experiments. All models demonstrate significantly better performance compared to the analogous experiment using standard prompting. This confirms that CoT is a highly effective technique for enhancing model efficacy. A notable phenomenon observed here is that almost all models exhibit high precision on the same-as relationship. However, this often comes at the cost of recall, with the highest recall achieved by openchat, reaching 0.780.
Among full open models, mixtral again outperformed the other model with an average F1-score of 0.808.
Regarding the quantised models, dolphin-mistral achieves the highest average F1-score of 0.869. It is closely followed by dolphin-mistral-dpo with a score of 0.826, solar at 0.825, and openchat at 0.823. This suggests that models fine-tuned on the OpenOrca corpus are particularly effective for this task. In this context, the use of the CoT strategy enables the quantised model dolphin-mistral to even surpass the much larger mixtral model.
Among the proprietary models, gpt-4 achieved the highest performance, reaching an average F1-score of 0.911. However, both gpt-4 and sonnet show a slight decrease in performance compared to other models. This suggests that some of the more advanced models might be confused rather than aided by the CoT strategy. On the other hand, both haiku and gpt-3.5 perform significantly better. Of course, without detailed knowledge of the pipelines of these more modern proprietary models, and whether they benefit from additional systems, it is difficult to draw definitive conclusions.
@@ 5.4 Chain-of-thought prompting, two-way strategy
The values are generally consistent with those obtained from the chain-of-thought one-way strategy experiments (see Table 3). Specifically, mixtral, dolphin-mistral, and gtp-4 emerge as the top-performing models among the full open models, the quantised models, and the proprietary models, respectively. The average F1 score increases by a few percentage points in 15 out of 17 models, confirming that the two-way strategy remains effective when combined with CoT.
@@ 5.5 Comparative Analysis
Table 5 presents the average F1-scores for all 17 models across the four strategies, including the relative differences between one-way and two-way strategies for both standard and chain-of-thought prompting. The results indicate that shifting from a one-way to a two-way strategy improved performance in 14 models for standard prompting and in 15 models for chain-of-thought prompting.
@@
Generally, we observe a gradual increase in the average F1-score from left to right in Table 5, demonstrating the beneficial effects of both CoT and the two-way strategy. Thirteen models showed overall improvement, while four models (llama-2, dolphin-openorca, gpt-4, and sonnet) experienced a decline in performance. When applying a 0.7 average F1-score as a quality benchmark, only dolphin-openorca and openchat surpassed this threshold under a standard prompting approach with a one-way strategy. However, with a two-way strategy in standard prompting, five models exceeded the threshold. This number increased to eight models when a chain-of-thought, one-way strategy was used. Ultimately, nine models achieved an average F1-score above the 0.7 threshold when employing a chain-of-thought, two-way strategy. These results underscore the critical importance of a good prompting strategy when working in a zero-shot setting.
@@
Finally, we can analyse the best model for each of the three categories: fully open models, quantised models, and proprietary models. The mixtral model stands out among fully open models, particularly when used with the standard prompting and two-way strategy (0.847 F1). In contrast, the dolphin-mistral model surpasses other quantised models when applied with the chain-of-thought prompting and two-way strategy (0.920 F1). Notably, dolphin-mistral outperforms mixtral, likely due to the additional fine-tuning on datasets like OpenOrca, which seem particularly beneficial for this task. Among proprietary models, sonnet delivers the best performance when using standard prompts and a straightforward, one-way approach (0.967 F1). Interestingly, unlike most open models, its performance declines with more complex prompts. Since sonnet is a black-box system that may incorporate additional components, pinpointing the exact cause of this behaviour is challenging.
| MODEL | Standard Prompting | CoT Prompting | ||||
| One-way | Two-way | Difference | One-way | Two-way | Difference | |
| mistral | 0.313 | 0.318 | 0.005 | 0.696 | 0.799 | 0.103 |
| mixtral | 0.779 | 0.847 | 0.068 | 0.808 | 0.843 | 0.035 |
| llama-2 | 0.669 | 0.605 | -0.064 | 0.576 | 0.534 | -0.042 |
| dolphin-mistral | 0.599 | 0.706 | 0.107 | 0.869 | 0.920 | 0.051 |
| dolphin-mistral-dpo | 0.552 | 0.795 | 0.243 | 0.826 | 0.906 | 0.080 |
| dolphin-openorca | 0.724 | 0.853 | 0.129 | 0.738 | 0.724 | -0.014 |
| openchat | 0.705 | 0.783 | 0.078 | 0.823 | 0.828 | 0.005 |
| openchat-gemma | 0.506 | 0.579 | 0.073 | 0.662 | 0.709 | 0.047 |
| solar | 0.646 | 0.710 | 0.064 | 0.825 | 0.860 | 0.035 |
| mistral-openorca | 0.613 | 0.630 | 0.017 | 0.739 | 0.772 | 0.033 |
| eurus | 0.466 | 0.488 | 0.022 | 0.706 | 0.739 | 0.033 |
| llama-3 | 0.568 | 0.562 | -0.006 | 0.605 | 0.649 | 0.044 |
| orca-2 | 0.372 | 0.491 | 0.119 | 0.700 | 0.720 | 0.020 |
| gpt-3.5 | 0.195 | 0.283 | 0.088 | 0.522 | 0.540 | 0.018 |
| gpt-4 | 0.948 | 0.962 | 0.014 | 0.911 | 0.933 | 0.022 |
| haiku | 0.712 | 0.794 | 0.082 | 0.806 | 0.839 | 0.033 |
| sonnet | 0.967 | 0.965 | -0.002 | 0.893 | 0.900 | 0.007 |
@@ 6 Error Analysis
The mixtral model misclassified 154 out of 1,000 topic pairs. Specifically, 75 pairs were incorrectly identified as broader, 63 as narrower, 16 as other, and none as same-as. mixtral demonstrates a highly conservative approach in predicting the same-as category, achieving a precision of 1, as indicated in Table 2. This comes at a cost, as the model exhibits a tendency to misclassify pairs that are labelled as same-as in the gold standard, often categorising them incorrectly as either broader or narrower. This misclassification results in a low recall for the same-as category (0.708), which represents a significant limitation of the model. Specifically, 29.2% of same-as pairs were misclassified: 16% were labelled as broader, 12.4% as narrower, and 0.8% as other. Additionally, the mixtral model occasionally confuses broader pairs with narrower ones and vice versa. Notably, 9.6% of broader pairs were incorrectly predicted, with 8% labelled as narrower and 1.6% as other. Similarly, 14.4% of narrower pairs were misclassified, with 10.4% labelled as broader and 4% as other.
The dolphin-mistral model exhibits a different pattern compared to mixtral, particularly excelling in pairs labelled as broader and narrower in the gold standard. It makes 80 misclassifications, predominantly in the same-as (39) and other (38) categories. Although the model still faces challenges with same-as relationships, it performs better than mixtral. Specifically, for the same-as relationships, a small percentage (6.8%) are misclassified as broader, while 8.8% are incorrectly identified as narrower.
@@ 7 Conclusions
We are currently working on several fronts. We are fine-tuning off-the-shelf models like BERT, Mistral, LLaMa 3, Flan T5, Gemma, and Phi 3 with the aim of enhancing the performance of lightweight models to reduce computational costs and environmental impact when generating ontologies of research areas. Furthermore, we plan to incorporate a semantic reasoner, such as RDFox, into our pipeline for identifying and resolving inconsistencies in the generated ontology. Finally, we aim to extend our analysis to other scientific fields such as Material Science, Medicine, and Physics, with the ultimate goal of creating a comprehensive, multi-field ontology of research topics.

### v2

@@ 5.2 Standard prompting, two-way strategy
Among the full open models, mixtral achieves an average F1-score of 0.847, significantly outperforming both llama-2 and mistral. It is able to identify the broader, narrower, and same-as with F1-score around 0.82, but shows exceptional capabilities (0.925) in determining the relationship other. The superior performance of mixtral in the two-way classification strategy likely arises from its mixture-of-experts architecture, which can dynamically activate different subsets of experts depending on the input [58]. Therefore, mixtral may be able to selectively engage experts specialized in reasoning along specific hierarchical directions [81].
Regarding the quantised models, there are three models (e.g., dolphin-mistral-dpo, dolphin-openorca, and openchat) that obtain an average F1-score above 0.78. The strong performance of these quantised models when employing the two-way strategy likely reflects recent advancements in their fine-tuning methodologies. Specifically, dolphin-mistral-dpo benefits from Direct Preference Optimization (DPO), a technique that improves the model’s capacity to follow complex instructions [82], while dolphin-openorca leverages demonstration-based learning from more capable teacher models [60].
@@
In this case, the full open models and their quantised counterparts exhibit similar performance levels (e.g., 0.847 for mixtral and 0.853 for dolphin-openorca). Moreover, the two-way strategy has effectively reduced the performance gap between these models and their larger, proprietary counterparts, narrowing the difference to just 0.113 F1 points. This enhancement suggests that the two-way strategy is effective in mitigating the performance gap between small and large models. This finding aligns with prior research demonstrating that similar self-verification mechanisms can improve model accuracy in information extraction tasks [83]. Indeed, the significant performance gains observed in open models indicate that they may be capable of accessing relevant knowledge when prompted within an appropriate verification framework.
@@ 5.3 Chain-of-thought prompting, one-way strategy
Table 3 presents the results of the CoT, one-way strategy experiments. All models demonstrate significantly better performance compared to the analogous experiment using standard prompting. This confirms that CoT is a highly effective technique for enhancing model efficacy. A notable phenomenon observed here is that almost all models exhibit high precision on the same-as relationship. However, this often comes at the cost of recall, with the highest recall achieved by openchat, reaching 0.780.
Among full open models, mixtral again outperformed the other model with an average F1-score of 0.808. As per Section 5.1, mixtral’s strong performance is likely attributed to its mixture-of-experts architecture and extensive pre-training on diverse datasets, including scientific and technical literature [58, 79].
Regarding the quantised models, dolphin-mistral achieves the highest average F1-score of 0.869. It is closely followed by dolphin-mistral-dpo with a score of 0.826, solar at 0.825, and openchat at 0.823. This suggests that models fine-tuned on the OpenOrca corpus are particularly effective for this task. The OpenOrca dataset’s inclusion of scientific and technical content likely contributes to this efficacy [60]. In this context, the use of the CoT strategy enables the quantised model dolphin-mistral to surpass the much larger mixtral model. This performance differential may stem from dolphin-mistral’s specialised fine-tuning on instruction-following datasets that emphasise reasoning processes [84].
Among the proprietary models, gpt-4 achieved the highest performance, reaching an average F1-score of 0.911. Notably, both gpt-4 and sonnet exhibit a slight decline in performance when using the CoT prompting strategy, whereas haiku and gpt-3.5 show improved results under the same conditions. The counterintuitive behavior of the larger models may arise from redundancy and cognitive overhead. Since these models have already internalized complex semantic relationships during pre-training, explicitly prompting them to articulate intermediate reasoning steps could introduce unnecessary noise [74, 85]. This additional complexity may disrupt their internal inference mechanisms, ultimately impairing performance and reducing accuracy.
@@ 5.4 Chain-of-thought prompting, two-way strategy
The values are generally consistent with those obtained from the CoT one-way strategy experiments (see Table 3). Specifically, mixtral, dolphin-mistral, and gtp-4 emerge as the top-performing models among the full open models, the quantised models, and the proprietary models, respectively. The average F1 score increases by a few percentage points in 15 out of 17 models, confirming that the two-way strategy remains effective when combined with CoT. The consistent improvement across models when employing the two-way strategy aligns with previous findings demonstrating that techniques aimed at enhancing logical consistency can lead to more robust judgments [76].
For quantised models, the good performance of dolphin-mistral further suggests that pretraining on OpenOrca [60] yields a consistent advantage on this task. In this case, the quantisation process [86] appears to preserve the model’s core representational capacity for handling technical terminology. This preservation of performance post-quantisation is particularly noteworthy and indicates that the model’s parameters encode knowledge efficiently, even under reduced precision constraints [87].
The variation in performance across different versions of the same base model (e.g., different mistral derivatives) suggests that fine-tuning procedures can either enhance or diminish domain-specific capabilities present in the base model [88]. This highlights the trade-offs inherent in optimising for general capabilities versus domain-specific knowledge retention [89]. The consistent improvement achieved through two-way verification with CoT reasoning across nearly all models suggests that, regardless of pre-training, explicit reasoning helps models more effectively leverage their knowledge of technical terminology and scientific concepts [90, 91].
@@ 5.5 Comparative Analysis
Table 5 presents the average F1-scores for all 17 models across the four strategies, including the relative differences between one-way and two-way strategies for both standard and CoT prompting. The results indicate that shifting from a one-way to a two-way strategy improved performance in 14 models for standard prompting and in 15 models for CoT prompting.
@@
Generally, we observe a gradual increase in the average F1-score from left to right in Table 5, demonstrating the beneficial effects of both CoT and the two-way strategy. Thirteen models showed overall improvement, while four models (llama-2, dolphin-openorca, gpt-4, and sonnet) experienced a decline in performance. When applying a 0.7 average F1-score as a quality benchmark, only dolphin-openorca and openchat surpassed this threshold under a standard prompting approach with a one-way strategy. However, with a two-way strategy in standard prompting, five models exceeded the threshold. This number increased to eight models when a CoT, one-way strategy was used. Ultimately, nine models achieved an average F1-score above the 0.7 threshold when employing a CoT, two-way strategy. These results underscore the critical importance of a good prompting strategy when working in a zero-shot setting.
@@
Finally, we can analyse the best model for each of the three categories: fully open models, quantised models, and proprietary models. The mixtral model stands out among fully open models, particularly when used with the standard prompting and two-way strategy (0.847 F1). In contrast, the dolphin-mistral model surpasses other quantised models when applied with the CoT prompting and two-way strategy (0.920 F1). Notably, dolphin-mistral outperforms mixtral, likely due to the additional fine-tuning on datasets like OpenOrca, which seem particularly beneficial for this task. Among proprietary models, sonnet delivers the best performance when using standard prompts and a straightforward, one-way approach (0.967 F1). Interestingly, unlike most open models, its performance declines with more complex prompts. Since sonnet is a black-box system that may incorporate additional components, pinpointing the exact cause of this behaviour is challenging.
| MODEL | Standard Prompting | CoT Prompting | Overall | ||||
| One-way | Two-way | Difference | One-way | Two-way | Difference | Improvement | |
| mistral | 0.313 | 0.318 | 0.005 | 0.696 | 0.799 | 0.103 | 0.486 |
| mixtral | 0.779 | 0.847 | 0.068 | 0.808 | 0.843 | 0.035 | 0.068 |
| llama-2 | 0.669 | 0.605 | -0.064 | 0.576 | 0.534 | -0.042 | 0.135 |
| dolphin-mistral | 0.599 | 0.706 | 0.107 | 0.869 | 0.920 | 0.051 | 0.321 |
| dolphin-mistral-dpo | 0.552 | 0.795 | 0.243 | 0.826 | 0.906 | 0.080 | 0.354 |
| dolphin-openorca | 0.724 | 0.853 | 0.129 | 0.738 | 0.724 | -0.014 | 0.129 |
| openchat | 0.705 | 0.783 | 0.078 | 0.823 | 0.828 | 0.005 | 0.123 |
| openchat-gemma | 0.506 | 0.579 | 0.073 | 0.662 | 0.709 | 0.047 | 0.203 |
| solar | 0.646 | 0.710 | 0.064 | 0.825 | 0.860 | 0.035 | 0.214 |
| mistral-openorca | 0.613 | 0.630 | 0.017 | 0.739 | 0.772 | 0.033 | 0.159 |
| eurus | 0.466 | 0.488 | 0.022 | 0.706 | 0.739 | 0.033 | 0.273 |
| llama-3 | 0.568 | 0.562 | -0.006 | 0.605 | 0.649 | 0.044 | 0.087 |
| orca-2 | 0.372 | 0.491 | 0.119 | 0.700 | 0.720 | 0.020 | 0.348 |
| gpt-3.5 | 0.195 | 0.283 | 0.088 | 0.522 | 0.540 | 0.018 | 0.345 |
| gpt-4 | 0.948 | 0.962 | 0.014 | 0.911 | 0.933 | 0.022 | 0.051 |
| haiku | 0.712 | 0.794 | 0.082 | 0.806 | 0.839 | 0.033 | 0.127 |
| sonnet | 0.967 | 0.965 | -0.002 | 0.893 | 0.900 | 0.007 | 0.074 |
@@ 6 Error Analysis
The mixtral model misclassified 154 out of 1,000 topic pairs. Specifically, 75 pairs were incorrectly identified as broader, 63 as narrower, 16 as other, and none as same-as. mixtral demonstrates a highly conservative approach in predicting the same-as category, achieving a precision of 1, as indicated in Table 2. This comes at a cost, as the model exhibits a tendency to misclassify pairs that are labelled as same-as in the gold standard, often categorising them incorrectly as either broader or narrower. This misclassification yields a low recall for the same-as category (0.708), which represents a significant limitation of the model. Specifically, 29.2% of same-as pairs were misclassified: 16% were labelled as broader, 12.4% as narrower, and 0.8% as other. Additionally, the mixtral model occasionally confuses broader pairs with narrower ones and vice versa. Notably, 9.6% of broader pairs were incorrectly predicted, with 8% labelled as narrower and 1.6% as other. Similarly, 14.4% of narrower pairs were misclassified, with 10.4% labelled as broader and 4% as other.
The dolphin-mistral model exhibits a different pattern compared to mixtral, particularly excelling in pairs labelled as broader and narrower in the gold standard. It makes 80 misclassifications, predominantly in the same-as (39) and other (38) categories. Although the model still faces challenges with same-as relationships, it performs better than mixtral. Specifically, for the same-as relationships, a small percentage (6.8%) are misclassified as broader, while 8.8% are incorrectly identified as narrower. Notably, the other relationships are misclassified equally as broader or narrower.
@@
7 Implications and limitations
The experiments presented in this study led to several notable insights. First, many models demonstrate strong zero-shot reasoning capabilities for this task, at least within the domain of engineering and computer science as represented by the IEEE Thesaurus. This performance may be partly due to the nature of relevant concepts in electrical engineering and computer science, which are likely well represented in the training data of LLMs, including code repositories, technical documentation, manuals, and patents. We intend to extend our analysis to other disciplines to determine whether this strong performance is consistent across fields and to evaluate the potential benefits of fine-tuning models for this specific task. Second, the performance of LLMs is strongly influenced by the choice of prompting strategy. In particular, CoT prompting clearly outperforms standard prompting, aligning with findings from prior research [74]. Additionally, running the classification process twice, switching the topic position, and applying simple rules to integrate the two outputs further reduces the number of errors. Indeed, the two-way CoT variant outperforms all other approaches, especially for smaller, more scalable, and cost-efficient models. The analysis indicates that selecting appropriate prompting strategies can lead to notable performance improvements, with gains exceeding 0.2 F1 points in several cases (see Table 5, last column). Third, smaller models achieve performance comparable to that of the best-performing architectures when adopting the best prompting strategy. Such models, particularly when fine-tuned on relevant datasets, can be highly effective for this task and may even outperform their base versions. For example, dolphin-mistral and dolphin-mistral-dpo surpass the standard mistral and are only outperformed by the largest proprietary models. These results suggest promising directions for the development of optimised and efficient models for ontology generation.
While we believe that these insights represent a substantial advancement in the state of the art, our current research settings present certain limitations that warrant discussion. We plan to address and explore these further in future work.
The first limitation lies in the domain-specific nature of our findings, which, as previously discussed, are currently confined to the IEEE taxonomy and the engineering field. A broader evaluation is required to assess the performance of LLMs in other domains. To this end, we are actively working on expanding our dataset to include classifications from additional disciplines, such as the Medical Subject Headings and the Physics Subject Headings, which represent the domains of biomedicine and physics, respectively. A second limitation is that our evaluation only focuses on LLM-based strategies. We do not compare these approaches with traditional AI methods, such as rule-based, statistical, or hybrid techniques. Nevertheless, we recognise the importance of such comparisons and intend to include them in future experiments. A third limitation arises from the dynamic nature of scientific research, where disciplines are continuously redefined by emerging paradigms and discoveries. Although our current results provide valuable insights, LLMs trained on present-day data may struggle to accurately classify relationships involving topics that have not yet emerged. Therefore, while our methodological approach offers long-term relevance, the specific LLMs employed will need to be updated periodically to reflect the most recent developments. We posit that future generations of LLMs will likely continue the observed trend of improvement, offering greater efficiency and enhanced performance at a reduced computational cost.
In addition, while this paper focused exclusively on identifying relationships among a subset of topics, future work will need to address the construction of a complete ontology. This entails resolving several challenges, including the evaluation of relationships across a much larger set of topics. To improve scalability and efficiency, it is crucial to identify strategies that reduce the search space. One promising approach proposed in the literature [92, 93] involves leveraging topic co-occurrence patterns in research paper abstracts and titles. Topic pairs that rarely or never co-occur can be automatically classified under the “other” category, thereby significantly reducing the computational cost of model inference.
Finally, while the results produced by LLMs demonstrate high accuracy, in realistic settings they must be further verified and refined by human experts. We envision a collaborative workflow in which LLMs generate initial ontology drafts that are subsequently reviewed and refined by domain experts. This process ensures both accuracy and the preservation of critical human judgment, which remains essential for developing robust and reliable KOS. Recent studies have proposed various strategies for integrating human-in-the-loop methodologies into the construction of knowledge graphs and ontologies [94].
@@ 8 Conclusions
We are currently working on several fronts. We are fine-tuning off-the-shelf models like BERT, Mistral, LLaMa 3, Flan T5, Gemma, Phi 3, and DeepSeek with the aim of enhancing the performance of lightweight models to reduce computational costs and environmental impact when generating ontologies of research areas. Furthermore, we plan to incorporate a semantic reasoner, such as RDFox [95], into our pipeline to identify and resolve inconsistencies. We will also make all the resulting ontologies publicly available, formatted according to the SKOS standard. Other directions include exploring explainability techniques, including SHAP values [96], attention visualization [97], and the identification of monosemantic features [98] to understand how LLMs process ontological relationships. We also plan to investigate potential biases in LLM-generated ontologies, such as examining whether inference error rates vary across different sub-domains of Engineering. Additionally, we will benchmark our findings against additional state-of-the-art ontology generation methods, including rule-based systems and statistical approaches. Ultimately, our goal is to develop a comprehensive, cross-disciplinary ontology of research topics. To this end, we plan to expand our analysis to cover a variety of scientific domains, including materials science, medicine, and physics.
@@
Acknowledgements
This publication is based upon work from COST Action CA23147 GOBLIN - Global Network on Large-Scale, Cross-domain and Multilingual Open Knowledge Graphs, supported by COST (European Cooperation in Science and Technology, https://www.cost.eu).